OpenStudy (anonymous):

What is the result of isolating y2 in the equation below? (x - 2)^2 + y2 = 64

OpenStudy (anonymous):

how would you isolate \(y_2\)?

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

i'm assuming it's\[\left( x-2 \right)^2+y_2=64\]

OpenStudy (anonymous):

couldn't you subtract off the bit that isn't \(y_2\)?

OpenStudy (anonymous):

these are my choices

OpenStudy (anonymous):

A. y2 = -x2 + 4x + 60 B. y2 = -x2 + 4x + 4 C. y2 = 64 - x2 D. y2 = x2 + 4x + 60

OpenStudy (anonymous):

it would be A?

OpenStudy (anonymous):

okay, but let's start by doing what makes sense and then work towards the form they give. what would you get if you subtracted (x-2)^2 from both sides and then expanded (multi[ied out) the (x-2)^2 and combined like terms?

OpenStudy (anonymous):

A is correct...good job!